Matt and Meg Comer are married and file a joint tax return. They do not have any children. Matt works as a history professor at a local university and earns a salary of $67,050. Meg works part-time at the same university. She earns $32,150 a year. The couple does not itemize deductions. Other than salary, the Comers’ only other source of income is from the disposition of various capital assets (mostly stocks). (Use the tax rate schedules ,Dividends and Capital Gains Tax Rates.) (Round final answers to the nearest whole dollar amount.) rev: 10_18_2018_QC_CS-144256 b. What is the Comers’ tax liability for 2018 if they report the following capital gains and losses for the year? Short-term capital gains $ 1,500 Short-term capital losses 0 Long-term capital gains 10,560 Long-term capital losses (10,230 )
Answer #1

Salary (67050+32150)

99200

Net short term capital gain (short term capital gain – short term capital loss = 1500-0=1500

1500

Net long term capital gain (long term capital gain – long term capital loss = 10560-10230 = 330

330

AGI

101030

Standard deduction (For MFJ 2018)

(24000)

Taxable income

77030

Qualified LTCG (preferentially taxed income)

(9000)

Income taxed at ordinary rates

68030

Ordinary Taxes

7783 (1905+(12%*(68030-19050)))

Preferential tax

0 (9000*0% as taxable income is below the limit of $77400)

Total tax liability

$7783 (ordinary taxes + preferential taxes)